OpenStudy (anonymous):

ILL GIVE YOU A MEDAL AND FAN!!! Write log3 6 as a logarithm of base 2.

OpenStudy (welshfella):

use the change of base formula for logs

OpenStudy (anonymous):

I dont know what that is.

OpenStudy (anonymous):

@texaschic101

OpenStudy (welshfella):

http://www.mathwords.com/c/change_of_base_formula.htm

Nnesha (nnesha):

change of base formula \[\huge\rm log_a b = \frac{ \log b }{ \log a }\]

Nnesha (nnesha):

\[\log = \log_{10}\]

OpenStudy (anonymous):

Im just kind of confused as to how I'm supposed to plug it in.. if that makes sense

OpenStudy (welshfella):

log2 6 = log3 6 ----- log3 2

OpenStudy (welshfella):

you can find log3 2 as a decimal y then log3 6 = y log2 6 = log2 6^y

OpenStudy (welshfella):

to find log3 2:- 2 = 3^y ln2 = y ln3 y = ln 2 / ln 3 = 0.6309 giving log3 6 = log2 6^(0.6309) = log2 3.097

OpenStudy (anonymous):

is that the answer?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

could you write it out as an equation? thats not one of the options...

OpenStudy (welshfella):

log3 6 = log2 6 / log2 3

OpenStudy (anonymous):

could you help me with this last one? Write log8 3 as a logarithm of base 5.

OpenStudy (welshfella):

iys very similar to the last one log8 3 = log5 3 / log5 8